OpenStudy (anonymous):

Write the given expression as an algebraic expression sin(arctan x- arcsin x)

OpenStudy (shubhamsrg):

sin(a-b) = .. ?

OpenStudy (anonymous):

\[put arc \tan x=\theta,x=\tan \theta \] \[arc \sin x=\Phi ,x=\sin \Phi \]  \[\sin \left( \theta-\Phi \right)=\sin \theta \cos \Phi-\cos \theta \sin \Phi \]\[\cos \Phi=\frac{ \sqrt{1-x ^{2}} }{1}=\sqrt{1-x ^{2}}\] \[\sin \theta=\frac{ x }{\sqrt{1+x ^{2}} },\cos \theta=\frac{ 1 }{ \sqrt{1+x ^{2}} }\] substitute the values

OpenStudy (anonymous):